# An Objective Improvement Approach

### v3

@@ 5.4. Making Games Sharp and Improving
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers777or from any other distribution over positive numbers that has 00 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
@@
We show the claim that, for two arbitrary but fixed (before drawing the $\alpha_{e}$) different valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ that satisfy the inequations in $H$ and any joint strategy $\sigma$, $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely.
As $\mathcal{G}$ is sharp, this implies for all joint strategies $\sigma$ and all neighbouring bases $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ that define a solution and the valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$, respectively, that they define that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely. (Note that sharpness implies $\mathsf{val}_{1}\neq\mathsf{val}_{2}$ in this case.) That is, it implies that $\mathcal{G}$ is improving.
@@ 5.5. Mixing Pivoting on the Simplex and of the Objective
Non-local updates are more complex than the other two, and the correctness proofs make use of the non-existence of the other two improvements. For both reasons, it seems natural to take non-local updates as a last resort.
@@ 6. Experimental Evaluation
Evaluation comes with a number of difficulties. First, there are no actual concrete benchmarks for DPGs. We have therefore translated parity games that model synthesis problems into DPGs. The translation process is exponential in the number of priorities of the parity games. For this reason we could translate and test only two types of concrete games: the Elevator and Language Inclusion. Besides these concrete games, we have also generated random games of various sizes. We have considered games with rational weights in the interval $[-250,250]$ on the edges, as games with integer weights represent a restricted class of games. The range of rational weights is not crucial as it suffice to multiply their values to scale the interval up or down. Moreover, we use a uniform discount factor of $0.9$, since for games in which each vertex has different weights on the edges, non-uniform discount factors do not make games any easier or harder to solve. Using a low discount factor (e.g. 0.5 or below) make, instead, games easier to solve, because with small discount factors the solution (i.e. the valuation $\mathsf{val}$) of the vertices converges after few steps, while with a discount factor close to 1 the valuation requires many steps to converge. Finally, it is worth noting that games with such random weights on the edges are also almost surely sharp, hence, it has not been necessarily to implement the sharpening technique described in Section 5.4.
@@
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 00000555510101010151515152020202025252525303030303535353540404040Game sizeNumber of iterations (avg)SIOI
@@
The number of iterations (Figure 2) represents how many times the linear programming solver has been called. As expected, SI is more efficient than OI at solving these games in terms of LP calls. What was less expected is how the advantage of SI over OI grows with the size of the game. Solving the linear programs, instead, is simpler for OI, at least when using a simplex style algorithm [^9] Simplex style algorithms (reported in Algorithm 1) operate in two phases: in the first phase, they look for any basis that defines a solution to the constraint system (Line 1), and in the second, they optimise (while loop). For OI, the first phase can be skipped for all but the first call; this is because the constraint system does not change, so that the previous solution can be used as a starting point. For SI, the constraint system changes; here, the previous solution is never a solution to the new linear program.. We also note that only 55 (5.5%) of the 1000 games considered overall reached a non-improving game. In these rare cases, this has been solved by choosing a stale switch (i.e. a switch of strategy that leads to the same valuation value).
@@
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 00045454545100100100100200200200200300300300300400400400400500500500500600600600600700700700700Game sizeNumber of updates (avg)SIOI
@@
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000001111222233334444555566667777888899991010101011111111Game sizeNumber of iterations (avg)SIOI
@@
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 0001601601601603003003003006006006006009009009009001,20012001{,}2001 , 2001,50015001{,}5001 , 5001,80018001{,}8001 , 8002,10021002{,}1002 , 100Game sizeNumber of updates (avg)SIOI
@@
Finally, for the concrete problems translated from parity games, all these games turned out to be easy to solve so that a single LP call suffices to find the optimal valuation. Therefore, in Table 2, instead of showing the number of iterations and updates, we show the solution time in seconds that provides a measure of the complexity of the game in terms of size of the linear system. Most of the games can be solved in less than one second. For the Elevator family of games, their size grows rapidly, so that we could not translate more than the first 5 instances. For the Language Inclusion games, instead, we report the games solved within 15 minutes.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000001111222233334444555566667777888899991010101011111111121212121313131314141414Game sizeNumber of iterations (avg)SIOI
@@
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , $0001{,}$ 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , 0005,50055005{,}5005 , 5006,00060006{,}0006 , 0006,50065006{,}5006 , 5007,00070007{,}0007 ,  [^0] Game sizeNumber of updates (avg)SIOI
@@
Elevator 1 $36$ $54$ 00
Elevator 2 $144$ $234$ 00
Elevator 3 $564$ $950$ 00
Elevator 4 $2688$ $4544$ $2$
@@
Language Inclusion 1 $170$ $1094$ 00
Language Inclusion 2 $304$ $1222$ 00
Language Inclusion 3 $428$ $878$ 00
Language Inclusion 4 $628$ $1538$ 00
Language Inclusion 5 $509$ $2126$ 00
Language Inclusion 6 $835$ $2914$ 00
Language Inclusion 7 $1658$ $4544$ $1$
@@ 7. Discussion
Seeing that (1) and (2) are equally cheap and simple, we could also favour (2) over (1): where we have a solution $\mathsf{val}$ defined by a basis, we can first find a joint strategy $\sigma$ such that $f_{\sigma}(\mathsf{val})$ is minimal among all strategies, before turning to (1) and (3). This would also make sure that no corner of the polytope is ever visited twice. This could happen in principle when favouring (1) over (2), as a step that is good for one objective function might not be good for its successor, but as the value of the objective function goes down in every step, a return to a previously optimal valuation is not possible.

### v4

@@ 5.4. Making Games Sharp and Improving
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers888or from any other distribution over positive numbers that has 0 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
@@
We show the claim that, for two arbitrary but fixed (before drawing the $\alpha_{e}$) different solutions $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ that satisfy the inequations in $H$ and any joint strategy $\sigma$, $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely.
@@
As $\mathcal{G}$ is sharp, this implies for all joint strategies $\sigma$ and all neighbouring bases $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ that define solutions $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$, respectively, that $\mathsf{val}_{1}\neq\mathsf{val}_{2}$ (by the sharpness of $\mathcal{G}$), which we showed implies that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely. That is, it implies that $\mathcal{G}$ is improving for the sampled set of positive numbers.
@@ 5.5. Mixing Pivoting on the Simplex and of the Objective
Non-local updates are more complex than the other two, and the correctness proofs make use of the optimality (w.r.t. $f_{\sigma}$) of the current solution. For both reasons, it seems natural to take non-local updates as a last resort.
@@ 6. Experimental Evaluation
Evaluation comes with a number of difficulties. First, there are no actual concrete benchmarks for DPGs. We have therefore translated parity games that model synthesis problems into DPGs. The translation process is exponential in the number of priorities of the parity games. For this reason we could translate and test only two types of concrete games: the Elevator and Language Inclusion. Besides these concrete games, we have also generated random games of various sizes. We have considered games with rational weights in the interval [250,250][-250,250] on the edges, as games with integer weights represent a restricted class of games. The range of rational weights is not crucial as it suffice to multiply their values to scale the interval up or down. Moreover, we use a uniform discount factor of $0.9$, since for games in which each vertex has different weights on the edges, non-uniform discount factors do not make games any easier or harder to solve. Using a low discount factor (e.g. 0.5 or below) make, instead, games easier to solve, because with small discount factors the solution (i.e. the valuation $\mathsf{val}_{\mathcal{G}}$) of the vertices converges after few steps, while with a discount factor close to 1 the valuation requires many steps to converge. Finally, it is worth noting that games with such random weights on the edges are also almost surely sharp, hence, it has not been necessarily to implement the sharpening technique described in Section 5.4.
@@
1001002002003003004004005005006006007007008008009009001,0001{,}0000551010151520202525303035354040Game sizeNumber of iterations (avg)SIOI
@@
The number of iterations (Figure 2) represents how many times the linear programming solver has been called. As expected, SI is more efficient than OI at solving these games in terms of LP calls. What was less expected is how the advantage of SI over OI grows with the size of the game. Solving the linear programs, instead, is simpler for OI, at least when using a simplex style algorithm [^10] Simplex style algorithms operate in two phases: in the first phase, they look for any basis that defines a solution to the constraint system (Line 1), and in the second, they optimise (while loop). For OI, the first phase can be skipped for all but the first call; this is because the constraint system does not change, so that the previous solution can be used as a starting point. For SI, the constraint system changes; here, the previous solution is never a solution to the new linear program.. We also note that only 55 (5.5%) of the 1000 games considered overall required a non-local improvement step. In these rare cases, this has been solved by choosing a stale switch (i.e. a switch of strategy that leads to the same solution).
@@
1001002002003003004004005005006006007007008008009009001,0001{,}0004545100100200200300300400400500500600600700700Game sizeNumber of updates (avg)SIOI
@@
1001002002003003004004005005006006007007008008009009001,0001{,}000011223344556677889910101111Game sizeNumber of iterations (avg)SIOI
@@
1001002002003003004004005005006006007007008008009009001,0001{,}0001601603003006006009009001,2001{,}2001,5001{,}5001,8001{,}8002,1002{,}100Game sizeNumber of updates (avg)SIOI
@@
Finally, for the concrete problems translated from parity games, all these games turned out to be easy to solve so that a single LP call suffices to find the optimal solution. Therefore, in Table 2, instead of showing the number of iterations and updates, we show the solution time in seconds that provides a measure of the complexity of the game in terms of size of the linear system. Most of the games can be solved in less than one second. For the Elevator family of games, their size grows rapidly, so that we could not translate more than the first 5 instances. For the Language Inclusion games, instead, we report the games solved within 15 minutes.
1001002002003003004004005005006006007007008008009009001,0001{,}000011223344556677889910101111121213131414Game sizeNumber of iterations (avg)SIOI
@@
1001002002003003004004005005006006007007008008009009001,0001{,}0001601605005001,0001{,}0001,5001{,}5002,0002{,}0002,5002{,}5003,0003{,}0003,5003{,}5004,0004{,}0004,5004{,}5005,0005{,}0005,5005{,}5006,0006{,}0006,5006{,}5007,0007{,} [^0] Game sizeNumber of updates (avg)SIOI
@@
Elevator 1 $36$ $54$ 0
Elevator 2 $144$ $234$ 0
Elevator 3 $564$ $950$ 0
Elevator 4 $2688$ $4544$ 22
@@
Language Inclusion 1 $170$ $1094$ 0
Language Inclusion 2 $304$ $1222$ 0
Language Inclusion 3 $428$ $878$ 0
Language Inclusion 4 $628$ $1538$ 0
Language Inclusion 5 $509$ $2126$ 0
Language Inclusion 6 $835$ $2914$ 0
Language Inclusion 7 $1658$ $4544$ 11
@@ 7. Discussion
Seeing that (1) and (2) are equally cheap and simple, we could also favour (2) over (1): where we have a solution $\mathsf{val}$ defined by a basis, we can first find a joint strategy $\sigma$ such that $f_{\sigma}(\mathsf{val})$ is minimal among all strategies, before turning to (1) and (3). This would also make sure that no corner of the polytope is ever visited twice. This could happen in principle when favouring (1) over (2), as a step that is good for one objective function might not be good for its successor, but as the value of the objective function goes down in every step, a return to a previously optimal solution is not possible.